Find the Probabilities on a standard normal curve


1. P(z>—1.32)

2. P(z< 0.45)

3. P(—2.76 < z < 2.76)

4. P(—3.0 < z < —1.68)

5. P(1.53 < z <2.91)


1
Expert's answer
2021-05-31T07:51:05-0400

1.) "P(Z>\u20141.32) = P(0<Z<1.32) = 0.90"


2.) "P(Z< 0.45) = 0.67"

3.) "P(\u20142.76 < Z < 2.76) = 2P(0<Z<2.76) = 0.99"


4.) "P(\u20143.0 < Z < \u20141.68) =P(0<Z<3)+P(0<Z<1.68) = 0.045"


5.) "P(1.53 < Z <2.91) = 0.061"